OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at (–5, 0) and a radius of 4?

hartnn (hartnn):

general form of circle with centre h,k and radius r is \(\large (x-h)^2+(y-k)^2 = r^2\)

hartnn (hartnn):

for your question h= -5 k=0 r =4 just plug in!

OpenStudy (anonymous):

(x + 5)^2 + (y - 0)^2 = 4^2

hartnn (hartnn):

correct! now simplify

OpenStudy (anonymous):

x^2 + 25 + y^2 = 16

OpenStudy (anonymous):

then I subtract 25 from both sides and get x^2 + y^2 = -9?

hartnn (hartnn):

\((x+5)^2 = x^2 +10x+25\)

OpenStudy (anonymous):

oh yeah so then (y - 0)^2 is just y^2?

hartnn (hartnn):

yes

hartnn (hartnn):

so, your equation is \(x^2+10x+y^2 = -9 \\ or, ~ x^2+y^2+10x+9=0\)

OpenStudy (anonymous):

(x + 5)^2 + y^2 = 4 x^2 + (y + 5)^2 = 4 (x + 5)^2 + y^2 = 16 x^2 + (y + 5)^2 = 16 are my choices.. so which one would it be?

OpenStudy (anonymous):

It's either A or C right?

hartnn (hartnn):

"(x + 5)^2 + (y - 0)^2 = 4^2" this was your reply , so \((x+5)^2 +y^2 =16\)

hartnn (hartnn):

its C because 4^2 =16

OpenStudy (anonymous):

okay thank you so much! I understand this now :)

hartnn (hartnn):

welcome ^_^